# Model-based Manipulation of Deformable Objects with Non-negligible Dynamics as Shape Regulation

### v1

@@ V-B Orientation
The experimental process is similar to that in Section V-A, now with $\zeta=(p_{e},\psi_{e,B^{\prime}})$ requiring the addition of the endpoint orientation error into the optimization cost. We simplify by ensuring the endpoint position goal $p^{*}$ is always exactly reachable by removing workspace position constraints on the manipulator, and only target endpoint orientation goals $\psi_{B^{\prime}}^{*}$ away from ±$\pm\pi$, so it is sufficient to define $d$ as the 2-norm of the combined position and angle errors. We relax the constraint on the manipulator angle to $\frac{-3\pi}{4}\leq\phi<\pi$, which allows a goal range of $\frac{-\pi}{2}\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$. Hence the optimization problem here is:
@@
Solving the problem for $\psi_{B^{\prime}}^{*}\in[\frac{-\pi}{2},\frac{\pi}{2}]$ at $\frac{\pi}{12}$ rad increments and $p^{*}=(0,0.4)$ produced the configurations shown in Fig. 11a for OB4. Fig. 11b displays a composite of the real object state at each of these configurations.
To quantitatively assess the accuracy, the angle of the tip was determined by measuring the points at the start and end of the endpoint weight to obtain a direction vector in the $xy$ plane. The error between this and the goal angle is plotted in Fig. 12. Along with the results for OB6, available here, this investigation also reveals considerable ability to reach specified orientations in the range $\frac{-\pi}{2}\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$, with a recorded maximum absolute angle error of around 0.29 rad or 17°. Particularly in the range of $0\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$ the error is maintained below 0.12 rad (7°) and 0.21 rad (12°) for OB4 and OB6 respectively.
Figure 11: a) Solutions for the endpoint orientation control experiment with OB4. b) Composite image showing real object configurations in the experiment. The cross indicates the intended endpoint position.
Figure 12: Measured tip angle error in the orientation control experiment with OB4.
@@ VI Conclusion
In this work, we have presented a novel approach to the manipulation of deformable linear objects using a strain-based modeling approach. The choice of a polynomial curvature parametrization allows for a compact manipulator-object system state that is suitable for model-based control. We then formally define the manipulation goal as a control problem, for which we introduce control strategies with provable guarantees. Our experimental results show that in this form, a model of minimal dimension is accurate enough for regulation of the steady-state endpoint position and orientation over a range of object properties. Our results demonstrate average positioning accuracy improvements of between 57.7% and 78.5% over a model-free reference approach and maximum orientation errors of 12°and 17°. We also show results assessing the accuracy of the full dynamic model. Future work will focus on closing the loop on the object state (non-collocated control) and performing more dynamic motion, starting from trajectory tracking and leading to accurate throwing and whipping.

### v2

@@ V-B Orientation
The experimental process is similar to that in Section V-A, now with $\zeta=(p_{e},\psi_{e,B^{\prime}})$ requiring the addition of the endpoint orientation error into the optimization cost. We simplify by ensuring the endpoint position goal $p^{*}$ is always exactly reachable by removing workspace position constraints on the manipulator, and only target endpoint orientation goals $\psi_{B^{\prime}}^{*}$ away from $\pm\pi$, so it is sufficient to define $d$ as the 2-norm of the combined position and angle errors. We relax the constraint on the manipulator angle to $\frac{-3\pi}{4}\leq\phi<\pi$, which allows a goal range of $\frac{-\pi}{2}\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$. Hence the optimization problem here is:
@@
Figure 17: Results from the object hooking task demonstration. The base position and object configurations generated for a series of endpoint goal locations are shown above, with corresponding images from the experiment below, overlaid with the intended object configuration at that time, and the radial constraint on the manipulator endpoint.
Figure 18: (a) The complete series of object configurations generated for the hooking task. (b) The measured trajectory of the object base during the experiment. Note that until $t=15s$ the manipulator is moving to the first generated trajectory point in the series. (c) The measured trajectory of the object endpoint coordinates, as well as the reference endpoint goal trajectory.
Solving the problem for $\psi_{B^{\prime}}^{*}\in[\frac{-\pi}{2},\frac{\pi}{2}]$ at $\frac{\pi}{12}$ rad increments and $p^{*}=(0,0.4)$ produced the configurations shown in the plots on the left of Fig. 16a, with composites of the real object state at each of these configurations displayed on the right.
To quantitatively assess the accuracy, the angle of the tip was determined by measuring the points at the start and end of the endpoint weight to obtain a direction vector in the $xy$ plane. The error between this and the goal angle is plotted in Fig. 16b. The position errors are also plotted in Fig. 16c. This investigation also reveals considerable ability to reach specified orientations in the range $\frac{-\pi}{2}\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$, with a recorded maximum absolute angle error of around 0.29 rad or 17°. Particularly in the range of $0\leq\psi_{B^{\prime}}^{*}\leq\frac{\pi}{2}$ the error is maintained below 0.12 rad (7°) and 0.21 rad (12°) for OB4 and OB6 respectively. The position errors are also small, with most cases being below $10\%$ of the object’s length.
V-C Task Demonstration
As a further demonstration using the controller in a task oriented manner, we provide an example where the robot needs to hang the end of the object over a hook which lies outside the reachable workspace of both the robot end effector, and the object endpoint when in the reference vertical configuration. Five main goal locations for the endpoint were defined to accomplish this task, with additional intermediate waypoints interpolated in between them. The first four locations involve varying the position of the object endpoint while keeping its orientation constant. The final location maintains a fixed position while rotating the end effector by $\pi/4$ radians. We then use the optimisation procedure for this series of goals to generate a trajectory for the object base pose, and execute this with the robot manipulator.
The upper row of Fig. 17 shows the generated configurations at the main goals, as well as the intermediate points, and images captured at these locations during the experiment are shown below, with the modeled object state and manipulator endpoint constraint overlaid (a video of the complete experiment is shown in the multimedia attachment). The complete set of configurations are shown together in Fig. 18a, and the measured trajectory of the object base and the x- and y-coordinates of the object endpoint along with its intended goal path are plotted in Fig. 18b and Fig. 18c, respectively.
@@ VI Conclusion
We presented a strain-based modeling approach for manipulating deformable linear objects, using a polynomial curvature parametrization developed within the context of soft robotics to enable compact, model-based control. The manipulation task was framed as a control problem, and controllers with provable guarantees were introduced. We focused in this work on planar bending (i.e., curvature only), but additional strains such as torsion can be added to demonstrate 3D manipulation. Results in simulations show that choosing a higher-order polynomial to describe the system’s curvature improves accuracy but comes at the cost of increased computational time when solving the optimization problems. Thus, for this paper, we opted for the linear curvature model as a compromise between accuracy and simplicity when it came to experimental validation. Experiments showed that even minimal models can achieve steady-state accuracy across various object properties, outperforming model-free baselines by 57.7–78.5% in positioning, with maximum orientation errors of 12°and 17°. We also demonstrated task-level applications and assessed dynamic model accuracy. Future work will focus on validating experimentally higher-order strain models, non-collocated feedback, dynamic motions involving trajectory tracking, and fast actions such as throwing or whipping.